OpenStudy (anonymous):

solve this rational equation x+ (10/x-2)=x^2+3x/x-2.. please! and this one 2/x+3+5/3-x=6/x^2-9 ..

OpenStudy (mathmale):

I'm going to try to interpret your question #1. Is the following what you meant?\[x+ (10/x-2)=x^2+3x/x-2\rightarrow?\rightarrow x+\frac{ 10 }{ x-2}=x^2+\frac{ 3x }{x-2 }?\]Wehn you wish to indicate division of 10 by (x-2), for example, it's very important that you use those parentheses. Otherwise, someone might interpret this as (10/x)-2. Verify your intent; then we can proceed.

OpenStudy (anonymous):

everything is right except the x^2 is suppose to be beside the 3x

OpenStudy (mathmale):

"beside the 3x" would result in your obtaining 3x*(x-2). I honestly don't think that that's what your posted homework problem calls for. Could you possibly use the Equation Editor or the Draw utility for greater clarity?

OpenStudy (anonymous):

x^2+3x over x-2

OpenStudy (mathmale):

Still doesn't look appropriate to me. Could you possibly do a screen shot and upload the screen shot to OpenStudy so that I could see the original question?

OpenStudy (mathmale):

x^2+3x over x-2 is exactly what I proposed earlier.

OpenStudy (anonymous):

ill send a picture

OpenStudy (mathmale):

\[ x+\frac{ 10 }{ x-2}=x^2+\frac{ 3x }{x-2}\] is what I THINK you mean.